# Class-Discriminative Attention Maps for Vision Transformers

### v2

@@ 3.4 Quantitative Evaluation
We perform quantitative evaluations of the correctness, class sensitivity, and compactness of importance estimators. In addition to our proposed methods (vanilla, smooth, and integrated CDAM), we perform evaluations on the raw attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], SmoothGrad [15], Integrated Gradient [16], and Input $\times$ Gradients. Note that based on classic LRP [28], partial LRP [22] provides improvement for multi-head self-attention mechanisms in modern transformer architecture [22]. Relevance propagation is a state-of-the-art importance estimator for explaining a prediction (e.g., class discriminatory) based on ViT [21]. We use 50 interpolation steps for Integrated Gradient and Integrated CDAM and chose the standard deviation of the noise used in SmoothGrad and smooth CDAM based on the best performance on the fidelity benchmark (explained below) on a held-out validation set of 200 images.
@@
To evaluate the correctness (also known as faithfulness) of importance scores, we employ two versions of input perturbation-based approaches [40, 41, 42]. Generally, an input feature (e.g., a pixel) is masked and a resulting model performance is observed. In all cases, input features are perturbed by being replaced with their blurred versions. A Gaussian blur with $\sigma=14$ turns pixels uninformative while minimizing perturbation artifacts [17].
@@
Fidelity. The first evaluation method uses accuracy curves obtained from the Most Important First (MIF) and the Least Important First (LIF) perturbations, which are abbreviated as MIF/LIF perturbation [17, 18]. Pixels are ranked by importance scores from a specific explainability method (e.g., CDAM) and are perturbed in the MIF or LIF order. In the MIF perturbation, an accurate importance score should lead to a relatively rapid decrease of the model output (logit) for the target class. For the LIF perturbation, we expect an inverse relationship. Note that some importance scores are unsigned (i.e., only non-negative values) and others are signed. For signed importance, with negative scores faithfully indicating counter-evidence, there may be an increase in the logits if pixels with negative scores are masked. For unsigned and correct importance scores, one expects initially constant logits in the LIF perturbation.
@@
The model output $f_{\text{MIF}}(p)$ or $f_{\text{LIF}}(p)$ as a function of the fraction of perturbed pixels $p$ gives the perturbation curve [17, 18]. We average these curves over all samples, providing an overview of fidelity for a given importance estimator. We further summarize the MIF and LIF perturbation evaluation by using the area $A_{\text{LIF-MIF}}$ (called the fidelity statistics) under the difference of perturbation curves $f_{\text{LIF}}-f_{\text{MIF}}$ to quantify the fidelity of the explanation methods [18]. Generally speaking, a higher value of fidelity statistics $A_{\text{LIF-MIF}}$ indicates a better explanation [17].
@@
Box sensitivity. For the second evaluation method, inspired by Sensitivity-n [43], we measure how the model output (logit) changes as a group of pixels are perturbed. Pixels are selected by a randomly placed square box (of size $s\times s$); then, the sum of their importance scores and the change in the model output are measured. For accurate importance estimators, we expect to see a high correlation between the two measures.
@@
In contrast to Sensitivity-n which randomly selects and perturbs $n$ pixels [43], we use a random square of pixels. A box removes a certain region which mimics how image patches are processed by ViT. Box sensitivity may therefore remove larger parts or complete objects. In contrast, removing random pixels may leave enough neighboring pixels with sufficient semantic information such that perturbation may not have any impact of the model output. Indeed, we have observed that for Sensitivity-n, the correlation is very low ($<0.1$) for all methods.
@@
measure the degree of class-discrimination. In particular, if an importance estimator is not sensitive to the target class at all, $\Delta f_{\text{LIF-MIF}}\approx\Delta f_{\text{box}}\approx 0$. An importance estimator that performs poorly in the data randomization test of [38] would have low $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$.
Compactness. We evaluate sparsity and shrinkage by counting the number of input pixels with importance scores smaller than $5\%$ of the maximum value. A threshold is defined by $t\times S_{c}^{\text{max}}$ with $t=0.05$, where $S_{c}^{\text{max}}$ is the maximum importance score for a given image and an importance estimator. Generally, a compact explanation is considered desirable, where sparsity and shrinkage are quantitative approaches [44].
@@ 3.5 Data
We perform quantitative evaluations on two datasets. The first one, $X_{\text{val}}$, is a randomly selected subset of the ImageNet validation set consisting of 1000 images [19]. The second one, $X_{\text{multi}}$, selects 1000 images from the validation set that contain multiple objects [20]. We filtered out instances in which the class labels provided by [20] actually refer to the same object (e.g. notebook and laptop) by calculating and thresholding the cosine similarity of the embedded class names. Embeddings were obtained using OpenAI’s text-embedding-3-large model.
@@
For application in medical images, we use lung computed tomography (CT) scans that were collected from and validated by seven academic centers and eight medical imaging companies [23]. The Lung Image Database Consortium image collection (LIDC) [23] contains 1018 records of CT scans. Presence of nodules (nodule $\leq$ 3 mm; nodule $<$ 3 mm; non-nodule $\geq$ 3 mm) is annotated by up to 4 human annotators. Segmentation of nodules (i.e., ROI) is the average contour given by annotators. Eight biomarkers that are informative in clinical practices are also collected in LIDC: subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter. For clinical descriptions of biomarkers, refer to [45]. After preprocessing, 854 nodule crops of size, corresponding biomarkers, and labels (benign or malignant) were saved and used for downstream procedures. Overall, the LIDC dataset used in this study contains 443 benign examples and 411 malignant examples.
@@ 4 Results
For our experiments using the ImageNet [19], we have been using the ViT-S/8 architecture trained with DINO as a backbone [9], which is available as a pre-trained model in their repository [^1] https://github.com/facebookresearch/dino. We first discuss the results of applying CDAM qualitatively and then perform a quantitative comparison with several baseline methods.
@@ 4.1.1 CDAM for a known class
To demonstrate our method, we have trained a linear classifier on the ImageNet [19] that achieved an accuracy of $77.0\%$. We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size of 128, and train for 10 epochs. The parameters of the ViT backbone are frozen during training, so only the classifier head is trainable. The predictions of this classifier were then used to create the heat maps shown in Fig. 1 via Eq. 2.
CDAM clearly distinguishes the region of interest (ROI) semantically related to the targeted class with positive values (yellow) from the rest of the image. Most backgrounds and other objects are assigned either zero (irrelevant) or negative scores (counter-evidence). Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in attention maps, see also the scatter plot between CDAM and attention map Fig. 11. CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class (Section A.1). This is a nice feature since CDAM inherits the high-quality object segmentations present in attention maps. This qualitative evaluation of shrinkage and sparsity of CDAM is reflected in the quantitative measures for compactness (Section 4.2.1).
One of the most useful features of CDAM is that it is highly sensitive to the chosen class. In Fig. 4 (left), when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable. Most of the other vegetables are assigned either near-zero or negative values (indicated in blue). Many other importance estimators do not distinguish the target class or only to a lesser degree. In Fig. 6, targeting the burger or the hotdog does not seem to change importance scores from partial LRP, Input $\times$ Grad, and IntGrad. Relevance propagation demonstrates some selective focus on the ROI, but is not as discriminative as CDAM (Fig. 6). This visual impression of class-sensitivity is supported by the quantitative evaluation in Section 4.2.2. Additional examples of CDAM and other interpretability methods are shown in the Appendix A.
Smooth CDAM appears to reduce the regularization of importance scores, highlighting, for instance, a larger portion of the truck and cars in the background in Fig. 5. This is quite different from how SmoothGrad behaves for CNNs [15], where it usually denoises the importance scores leading to a stronger focus on the ROI. On the other hand, Integrated CDAM can’t, for most samples, be visually distinguished from vanilla CADM even though its performance in the benchmarks is quite different (below).
@@ 4.1.2 CDAM for a latent concept
Concept embeddings $l_{c}$ have been obtained by averaging 30 latent representations of images depicting the concept of interest. These 30 images were randomly selected from the corresponding classes in the ImageNet validation set [19] and the CDAM in Fig. 3 were obtained using Eq. 4. We observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image. For example, for the concept hammer (bottom right), CDAM primarily highlights the hammer. CDAM also provides positive importance scores for the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept. Unrelated tools or parts appear in blue indicating that the model deems them to be unrelated to the concept hammer.
Just as CDAM is sensitive to the chosen class, it also produces attention maps that are sensitive to the chosen concept (Fig. 4, right). This makes the proposed method particularly valuable for understanding the representations the ViT has learned, since the explanations for different concepts can give a complementary view. For instance, CDAM distinguishes african elephant and zebra. However, CDAM also shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig. 4, right) which is initially counter-intuitive. However, when targeting the concept zebra, the model deems this part of the trunk to be similar to a zebra. Speculatively, the trunk may be too similar to the zebra’s muzzle or the trunk’s wrinkles may be mistaken for the zebra’s stripes.
@@ 4.2.2 Evaluation of class discrimination
The importance scores for CDAM become anti-correlated with the model output when the wrong importance scores are used for the perturbation (Fig. 10). This anti-correlation probably results from the fact that pixels that are evidence for the correct target class are counter-evidence for the wrong class. This result corroborates the visual impression that CDAM correctly assigns importance scores with opposite signs to the objects corresponding to the targeted class and other objects that are present (Fig. 6). This is also reflected in the negative values of $f_{\text{LIF}^{\prime}}-f_{\text{MIF}^{\prime}}$ for CDAM (Fig. 9), indicating that the ranking from least to most important has been, at least partially, reversed.

### v3

@@ 3.4 Quantitative Evaluation
We perform quantitative evaluations of the correctness, class sensitivity, and compactness of importance estimators. In addition to our proposed methods (vanilla, Smooth, and Integrated CDAM), we perform evaluations on attention maps, Input $\times$ Gradients, relevance propagation (RP) (Chefer et al., 2021a), partial Layer-wise Relevance Propagation (LRP) (Voita et al., 2019), SmoothGrad (Smilkov et al., 2017), IntGrad (Sundararajan et al., 2017), and GradCam for ViT (Gildenblat, 2021). Note that based on classic LRP (Bach et al., 2015), partial LRP (Voita et al., 2019) provides improvement for multi-head self-attention mechanisms in modern transformer architecture (Voita et al., 2019). Relevance propagation is a state-of-the-art importance estimator for explaining a prediction (e.g., class discriminatory) based on ViT (Chefer et al., 2021a). For IntGrad, SmoothGrad, Integrated CDAM, and Smooth CDAM, $n=50$ is used. The noise ($\sigma$) used in SmoothGrad and Smooth CDAM is selected based on the best performance on a validation set of 200 images.
@@
To evaluate the correctness (also known as faithfulness) of importance scores, we employ two versions of input perturbation-based approaches (Samek et al., 2016; Petsiuk et al., 2018; Kindermans et al., 2017). Generally, an input feature (e.g., a pixel) is masked and a resulting model performance is observed. In all cases, input features are perturbed by being replaced with their blurred versions. A Gaussian blur with $\sigma=14$ turns pixels uninformative while minimizing perturbation artifacts (Brocki & Chung, 2023c).
@@
Fidelity. The first evaluation method uses accuracy curves obtained from the Most Important First (MIF) and the Least Important First (LIF) perturbations, which are abbreviated as MIF/LIF perturbation (Brocki & Chung, 2023c; b). Pixels are ranked by importance scores from a specific explainability method (e.g., CDAM) and are perturbed in the MIF or LIF order. In the MIF perturbation, an accurate importance score should lead to a relatively rapid decrease of the model output (logit) for the target class. For the LIF perturbation, we expect an inverse relationship. Note that some importance scores are unsigned (i.e., only non-negative values) and others are signed. For signed importance, with negative scores faithfully indicating counter-evidence, there may be an increase in the logits if pixels with negative scores are masked. For unsigned and correct importance scores, one expects initially constant logits in the LIF perturbation.
@@
The model output $f_{\text{MIF}}(p)$ or $f_{\text{LIF}}(p)$ as a function of the fraction of perturbed pixels $p$ gives the perturbation curve (Brocki & Chung, 2023c; b). We average these curves over all samples, providing an overview of fidelity for a given importance estimator. We further summarize the MIF and LIF perturbation evaluation by using the area $A_{\text{LIF-MIF}}$ (called the fidelity statistics) under the difference of perturbation curves $f_{\text{LIF}}-f_{\text{MIF}}$ to quantify the fidelity of the explanation methods (Brocki & Chung, 2023b). Generally speaking, a higher value of fidelity statistics $A_{\text{LIF-MIF}}$ indicates a better explanation (Brocki & Chung, 2023c). Similar ideas are developed into the symmetric relevance gain (SRG) measure (Bluecher et al., 2024).
@@
Box sensitivity. For the second evaluation method, inspired by Sensitivity-n (Ancona et al., 2017), we measure how the model output (logit) changes as a group of pixels is perturbed. Pixels are selected by a randomly placed square box (of size $s\times s$); then, the sum of their importance scores and the change in the model output are measured. For accurate importance estimators, we expect to see a high correlation between the two measures.
@@
In contrast to Sensitivity-n which randomly selects and perturbs $n$ pixels (Ancona et al., 2017), we use a random square of pixels. A box removes a certain region which mimics how image patches are processed by ViT. Box sensitivity may therefore remove larger parts or complete objects. In contrast, removing random pixels may leave enough neighboring pixels with sufficient semantic information such that perturbation may not have any impact of the model output. Indeed, we have observed that for Sensitivity-n, the correlation is very low ($<0.1$) for all methods.
@@
measure the degree of class-discrimination. In particular, if an importance estimator is not sensitive to the target class at all, $\Delta f_{\text{LIF-MIF}}\approx\Delta f_{\text{box}}\approx 0$. An importance estimator that performs poorly in the data randomization test of (Adebayo et al., 2018) would have low $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$.
Compactness. We evaluate sparsity and shrinkage by counting the number of input pixels with importance scores smaller than $5\%$ of the maximum value. A threshold is defined by $t\times S_{c}^{\text{max}}$ with $t=0.05$, where $S_{c}^{\text{max}}$ is the maximum importance score for a given image and an importance estimator. Generally, a compact explanation is considered desirable, where sparsity and shrinkage are quantitative approaches (Nauta et al., 2023a).
@@ 3.5 Data
We have used two subsets of the ImageNet (denoted as $X_{\text{val}}$ and $X_{\text{multi}}$) (Deng et al., 2009), as well as lung computed tomography (CT) scans from the Lung Image Database Consortium image collection (LIDC) (Armato III et al., 2011). The first one, $X_{\text{val}}$, is a randomly selected subset of the ImageNet validation set consisting of 1000 images (Deng et al., 2009). The second one, $X_{\text{multi}}$, selects 1000 images from the validation set that contain multiple objects (Beyer et al., 2020). We filtered out instances in which the class labels provided by (Beyer et al., 2020) actually refer to the same object (e.g. notebook and laptop) by calculating and thresholding the cosine similarity of the embedded class names. Embeddings were obtained using OpenAI’s text-embedding-3-large model.
@@
The Lung Image Database Consortium image collection (LIDC) contains 1018 records of lung CT scans, that were collected from and validated by seven academic centers and eight medical imaging companies (Armato III et al., 2011). Presence of nodules (nodule $\leq$ 3 mm; nodule $<$ 3 mm; non-nodule $\geq$ 3 mm) is annotated by up to 4 human annotators. Segmentation of nodules (i.e., ROI) is the average contour given by annotators. Eight biomarkers that are informative in clinical practices are also collected in LIDC: subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter. For clinical descriptions of biomarkers, refer to (Opulencia et al., 2011). After preprocessing, 854 nodule crops of size, corresponding biomarkers, and labels (benign or malignant) were saved and used for downstream procedures. Overall, the LIDC dataset used in this study contains 443 benign examples and 411 malignant examples.
@@ 4 Results
For our experiments using the ImageNet (Deng et al., 2009), we use the ViT-S/8 architecture (with a patch size of $8\times 8$) trained with DINO as a backbone (Caron et al., 2021). Instead of using a ViT model that is pre-trained in a self-supervised manner, we demonstrate CDAMs directly from an alternative ViT architecture trained with supervised weakly through hashtags (SWAG) (Singh et al., 2022) in Section A.5. We further demonstrate applicability of CDAM on computed tomography (CT) scans, using DINO (Caron et al., 2021) in Section 5, as well as using Data-efficient image Transformers (DeIT) (Touvron et al., 2021) and DINOv2 (Oquab et al., 2024) in Section A.5.2.
@@ 4.1.1 CDAM for a known class
Figure 3: Visual comparison of different importance estimators which obtain explanations w.r.t. different output classes. Similar heat maps regardless of target classes imply that importance estimators are not class discriminative. Attention maps (AM) are, by design, not class discriminative. Orange and blue colors correspond to positive and negative values, respectively.
To demonstrate our method, we have trained a linear classifier on the ImageNet (Deng et al., 2009) that achieved an accuracy of $77.0\%$. We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size of 128, and train for 10 epochs. The parameters of the ViT backbone are frozen during training, so only the classifier head is trainable. The predictions of this classifier were then used to create the heat maps shown in Fig. 1 via Eq. 2.
CDAM clearly distinguishes the region of interest (ROI) semantically related to the targeted class with positive values (yellow) from the rest of the image. Most backgrounds and other objects are assigned either zero (irrelevant) or negative scores (counter-evidence). Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in attention maps, see also the scatter plot between CDAM and attention map Fig. S1. CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class. We discuss the proportionality of CDAM to the attention scores in Section A.1. It is a nice feature since CDAM inherits the high-quality object segmentations present in attention maps. This qualitative evaluation of shrinkage and sparsity of CDAM is reflected in the quantitative measures for compactness (Section 4.2.1).
One of the most useful features of CDAM is that it is highly sensitive to the chosen class. In Fig. 5 (left), when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable. Most of the other vegetables are assigned either near-zero or negative values (indicated in blue). Many other importance estimators do not distinguish the target class or only to a lesser degree. In Fig. 3, targeting the burger or the hotdog does not seem to change importance scores from partial LRP, Input $\times$ Grad, and IntGrad. Relevance propagation demonstrates some selective focus on the ROI, but is not as discriminative as CDAM (Fig. 3). This visual impression of class-sensitivity is supported by the quantitative evaluation in Section 4.2.2. Additional examples of CDAM and other interpretability methods are shown in Section A.3.
@@ 4.1.2 CDAM for a latent concept
Figure 4: Class-Discriminative Attention Maps (CDAM) for user-defined concepts (Section 3.2). The concept embedding $l_{c}$ has been obtained by averaging the latent representations of 30 images that include the common concept. In each instance, from left to right: Sample image, AM, and CDAM. Orange and blue correspond to positive and negative values, respectively.
Concept embeddings $l_{c}$ have been obtained by averaging latent representations of 30 images. Those images were randomly selected from a class in the ImageNet validation set (Deng et al., 2009). Then, CDAMs in Fig. 4 were obtained using Eq. 4. We observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the latent concepts shared by selected images from the rest of the image. For example, when having selected images containing hammer (bottom right), CDAM primarily highlights the hammer. While the annotated classes were used for this experiment, the concepts shared by those images are likely more nuanced, specific, and complex than the class itself. For example, CDAM also provides positive importance scores for the wooden handles of the screwdrivers and the socket attached the ratchet, which imply relevance for the concepts shared by 30 images selected from the class hammer. Semantically, 30 images that are selected from a class hammer have a wooden handle (i.e., concept), which is shared by multiple tools. Similarly, the socket at the top of the ratchet shares similarities with typical metal heads at the top of the hammer. Unrelated tools or parts appear in blue indicating that the model deems them to be unrelated to the concept hammer.
Figure 5: When targeting different classes (left) or concepts (right), the resulting CDAMs are distinct and align with the corresponding objects. Note that green parts of vegetables are shown to activate the zuccini class (left). The wrinkles in the elephant’s trunk appear to be related to the zebra concept, semantically mistaken by the model as zebra stripes.
CDAMs are discriminative with respect to the target concepts shared by selected samples (Fig. 5, right). This makes the proposed method particularly valuable for understanding the representations the ViT has learned, since the explanations for different concepts can give a complementary view. For example, by selecting 30 images that contain african elephant, CDAM highlights areas that are important for concepts which coincide largely with the body of an elephant. It also shows that parts of the elephant’s trunk are counter-evidence for the concepts shared by images with african elephant (Fig. 5, right). For the same sample image, we can also select 30 images with zebra, which suggests that the model deems the part of the elephant’s trunk to be similar to concepts shared by those images.
Although we selected a group of images based on annotated classes, which are denoted in Fig. 5 (right), to illustrate concept-based CDAMs, generally the concepts shared by selected images are not identical to the classes themselves. One could manually provide a set of images with a shared concept that is not originally annotated in the dataset. For example, 10 images that contain zebras may share semantic concepts (e.g., “stripes” and “tail”), that make up the class zebra. We demonstrate this by selecting a set of 10 images, outside of the ImageNet, that exclusively show stripes and another set of 10 images that contain diverse animal tails. Shown in Fig. 6, CDAM for stripes highlights the body of the zebra, but not the body of the cheetah. On the other hand, CDAM for tails tends to focus on the tail of the cheetah. Note that 10 images of diverse animal tails include body parts of animals, often rear ends and legs.
Figure 6: Concept-based CDAM based on 10 images that show stripes or contain animal tails.
@@ 4.2.1 Evaluation of correctness
Fidelity Box Sensitivity
[…]
CDAM 404 1038 632 12.2
Integrated CDAM 375 1045 669 13.0
Smooth CDAM 282 1110 827 12.4
Relevance Propagation 315 957 643 10.0
Partial LRP 489 837 348 4.5
Attention map 452 846 393 4.8
SmoothGrad 534 529 -4 0.4
IntGrad 226 802 574 7.6
GradCAM 300 944 643 8.6
Input $\times$ Grad 273 713 439 3.5
Table 1: Evaluation of correctness by fidelity and box sensitivity. In $X_{\text{multi}}$, importance scores for the target class were used to perturb pixels. The changes in the model output w.r.t. the perturbation percentage ($f_{\text{LIF-MIF}}$; Fig. S2) or the box size ($f_{\text{box}}$; Fig. S3) are quantified by the area under the curves.
@@ 4.2.2 Evaluation of class discrimination
Fidelity Box Sensitivity
[…]
CDAM 739 14.6
Integrated CDAM 708 14.8
Smooth CDAM 823 13.4
Relevance Propagation 293 6.6
Partial LRP 3 0.0
Attention map 0 -0.2
SmoothGrad -11 -0.1
IntGrad 540 7.6
GradCAM 681 9.4
Input $\times$ Grad 433 3.4
Table 2: Evaluation of class discrimination by differences in fidelity and box sensitivity metrics. By using the $X_{\text{multi}}$ dataset, we measure how sensitive importance scores are to changing their target class (Details in Section 3.5). Area under the $\Delta f_{\text{LIF-MIF}}$ (Fig. S4) and $\Delta f_{\text{box}}$ curves (Fig. S5).
@@
The importance scores for CDAM become anti-correlated with the model output when the wrong importance scores are used for the perturbation (Fig. S5). This anti-correlation probably results from the fact that pixels that are evidence for the correct target class are counter-evidence for the wrong class. This result corroborates the visual impression that CDAM correctly assigns importance scores with opposite signs to the objects corresponding to the targeted class and other objects that are present (Fig. 3). This is also reflected in the negative values of $f_{\text{LIF}^{\prime}}-f_{\text{MIF}^{\prime}}$ for CDAM (Fig. S4), indicating that the ranking from least to most important has been, at least partially, reversed.